# LinVT: Empower Your Image-level Large Language Model to Understand Videos

### v1

@@ 3.2.1 Spatio-Temporal Visual Token Refiner (SVR)
For the visual token $V_{i}\in\mathbb{R}^{H\times W\times C}$ of the $i$-frame, where $H$, $W$ and $C$ represent the height, width and the number of channels, respectively. The spatial significance score for each frame is obtained by applying self-attention within the corresponding visual token. Subsequently, the final significance scores are computed by applying self-attention across the sequence of spatial significance scores for the frames, thereby incorporating temporal information.
@@
In this work, we conduct a rigorous investigation of the designs of multi-scale token processing. Specifically, three variants are compared, as illustrated in Fig. 3 (right part). Multi-A: MTP is placed after TTA. In this case, pooling is applied on the video tokens generated by TTA (see details of TTA in Sec. 3.2.2). Multi-B and Multi-C: MTP is placed before TTA, and MTP operates on the top-$k$ selected visual tokens. The only difference between these two designs is whether TTA is applied on all scales at once (i.e. Multi-B) or each scale separately (i.e. Multi-C). By comparing these three designs experimentally (see Sec. 4.1), Multi-C is the chosen design in our LinVT. The output of MTP is a set of multi-scale video tokens $MT=\{T^{1},T^{2},\ldots,T^{l}\}$, where $l$ is the number of scales.
@@ 3.2.2 Text-conditioned Token Aggregation (TTA)
The TTA module is designed to refine and integrate multi-scale visual tokens with textual information, thereby enhancing the model’s capacity to understand and interpret video content within a multimodal context. To achieve this, scale-specific queries are utilized to aggregate multi-scale visual tokens, which not only reduce the number of visual tokens but also integrate and enhance the semantic information at a higher level. Furthermore, we have developed a cross-modal decoder to combine image and text modal features. Each group of cross-modal queries is processed through a self-attention layer, an image cross-attention layer (to integrate image features), and a text cross-attention layer (to integrate text features). This approach injects textual information into the queries to achieve improved modal alignment.
Scale-specific learnable queries. In TTA, we utilize scale-specific learnable queries for dynamic information extraction from the visual tokens. Compared to previous methods [1, 29] that process all tokens equally, utilizing scale-specific queries is a novel approach that allows the model to focus on the information across different scales. The queries are learnable, which are represented as $MQ=\{\mathbf{Q}^{1},\mathbf{Q}^{2},\ldots,\mathbf{Q}^{l}\}$, where $l$ is the number of pre-defined scales. Importantly, the queries in each scale only interact with the visual tokens of that particular scale. Therefore, the TTA module operates $l$ times in total.
Visual-query interaction. After the self-attention process, the temporal attention queries and visual tokens from the SVR module engage in cross-attention. This interaction helps to refines these visual tokens by focusing on specific aspects of the visual content that are relevant to the task.
@@
Discussion. The design of our TTA module differs from existing methods in several ways. First, in comparison to Q-Former [11], TTA takes a more targeted approach in terms of multi-scale aggregation and text-driven visual token extraction, particularly in the context of video understanding. Our design places greater emphasis on handling temporal information and fusing multi-modal information. Moreover, while Perceiver [1] also utilizes cross-attention mechanisms, the TTA module specifically adapts the query structure to accommodate temporal multi-scale characteristics. This further enhances the role of text in visual token compression, enabling TTA to effectively integrate temporal and semantic information in video understanding tasks.
@@ 4.2 Comparison with Sate-of-The-Art
Table 8: Comparison with the state-of-the-art on six shot and general video benchmarks. The best results are in bold and second best underlined.

### v2

@@ 3.2.1 Spatio-Temporal Visual Token Refiner (SVR)
Concretely, for the visual token $V_{i}\in\mathbb{R}^{H\times W\times C}$ of the $i$-frame, where $H$, $W$ and $C$ represent the height, width and the number of channels, respectively. The spatial significance score for each frame is obtained by applying self-attention within the corresponding visual token. Subsequently, the final significance scores are computed by applying self-attention across the sequence of spatial significance scores for the frames, thereby incorporating temporal information.
@@
In this work, we conduct a rigorous investigation of the designs of multi-scale token processing. Specifically, three variants are compared, as illustrated in Fig. 3 (right part). Multi-A: MTP is placed after TTA. In this case, pooling is applied on the video tokens generated by TTA (see details of TTA in Sec. 3.2.2). Multi-B and Multi-C: MTP is placed before TTA, and MTP operates on the top-$k$ selected visual tokens. The only difference between these two designs is whether TTA is applied on all scales at once (i.e. Multi-B) or each scale separately (i.e. Multi-C). By comparing these three designs experimentally (see Sec. 4.1), Multi-C is the chosen design in our LinVT. The output of MTP is a set of multi-scale video tokens $MT_{v}=\{T_{v}^{1},T_{v}^{2},\ldots,T_{v}^{l}\}$, where $l$ is the number of scales.
@@ 3.2.2 Text-conditioned Token Aggregation (TTA)
The TTA module is designed to refine and integrate multi-scale visual tokens with textual information, thereby enhancing the model’s capacity to understand and interpret video content within a multimodal context. To achieve this, scale-specific queries are utilized to aggregate multi-scale visual tokens, which not only reduce the number of visual tokens but also integrate and enhance the semantic information at a higher level. Furthermore, we have developed a cross-modal decoder to combine image and text modal features. Each group of cross-modal queries is processed through a self-attention layer, an image cross-attention layer (to integrate image features), and a text cross-attention layer (to integrate text features). Lastly, the processed cross-modal queries are fed into the linear aggregation layer to perform linear aggregation on the multi-scale visual tokens. This approach injects textual information into the queries to achieve improved modal alignment.
Scale-specific learnable queries. In TTA, we utilize scale-specific learnable queries for dynamic information extraction from the visual tokens and text tokens. Notably, the learnable queries are designed to generate attention maps across the visual tokens, which takes place in the final linear aggregation layer. Compared to previous methods [1, 29] that process all tokens equally, utilizing scale-specific queries is a novel approach that allows the model to focus on the information across different scales. The queries are learnable, which are represented as $MQ=\{\mathbf{Q}^{1},\mathbf{Q}^{2},\ldots,\mathbf{Q}^{l}\}$, where $l$ is the number of pre-defined scales. The self-attention mechanism operates exclusively within the same scales, without any attention across different scales.
Visual-query interaction. After the self-attention process, the temporal attention queries and visual tokens from the SVR module engage in cross-attention. Note that the queries in each scale only interact with the visual tokens of the corresponding scale. This interaction helps to refines these visual tokens by focusing on specific aspects of the visual content that are relevant to the task.
@@
Linear aggregation layer. The linear aggregation layer begins by calculating attention maps using the scale-specific queries $Q_{v,t}$ (after $L$ iterations of self-attention, visual-query and text-query interactions), in conjunction with the multi-scale visual tokens ($T_{v}$). These attention maps are subsequently used to aggregate the multi-scale visual tokens into fixed-length tokens. In other words, the output tokens of TTA represent various configurations of the weighted sum of multi-scale visual tokens, thereby enforcing linearity:
$\Phi_{Lin}(T_{v},Q_{v,t})$ $=W_{attn}\cdot T_{v}$
$=softmax(W_{q}\cdot Q_{v,t}~{},~{}W_{k}\cdot T_{v})\cdot T_{v},$
where $\Phi_{Lin}$ is a linear function that computes different weighted sum of the visual tokens $T_{v}$, and $W_{attn}$ is the attention map derived from $T_{v}$ and ${Q_{v,t}}$. In contrast to vanilla cross-attention, $\Phi_{Lin}$ does not include residual connections and directly utilizes input values $T_{v}$ without any transformations.
Discussion. The design of our TTA module differs from existing methods like Q-Former [11] and Perceiver [1] in several ways. First, the output of TTA is a linear combination of its input. Second, TTA takes a more targeted approach in terms of multi-scale aggregation and text-driven visual token extraction, particularly in the context of video understanding. Our design places greater emphasis on handling temporal information and fusing multi-modal information. Moreover, while Perceiver also utilizes cross-attention mechanisms, the TTA module specifically adapts the query structure to accommodate temporal multi-scale characteristics. This further enhances the role of text in visual token compression, enabling TTA to effectively integrate temporal and semantic information in video understanding tasks.
@@ 4.2 Comparison with Sate-of-The-Art
Table 8: Comparison with the state-of-the-art on six short and general video benchmarks. The best results are in bold and second best underlined.